OpenStudy (anonymous):

how do i find the maximum angle to keep a 12kg block stationary on an incline plane?

OpenStudy (zephyr141):

i guess you would need to know the coefficient of static friction to answer this.

OpenStudy (amistre64):

lift the plane until the block starts to slide .....

OpenStudy (amistre64):

in physics class we did something like this, forgot what it was we were trying to find tho .... prolly one of the friction coeffs

OpenStudy (zephyr141):

\[\sum F_{y}:n=mgcos \theta \]\[\sum F_{x}:-f_{s}+mgsin \theta =ma\]\[\mu_{s}=\frac{1-\sin\theta}{\cos\theta}\]i'm just trying to solve this... i'm not sure i did it even remotely correct but there's something. just trying to isolate theta. \[\mu_{s}=\frac{-1}{\cos\theta}+\frac{\sin\theta}{\cos\theta}\]\[\mu_{s}=-\csc\theta+\tan\theta\]so yeah i'm just throwing stuff down now.... help someone.